OpenStudy (nightmarenight):

What rotation symmetry does a square have? 30 degrees 45 degrees 90 degrees 60 degrees

OpenStudy (anonymous):

@jim_thompson5910 gave you a great defintion... "For any regular n-gon, it has a rotational symmetry of 360/n For example, if n = 5, we have a pentagon. If we have a regular pentagon, then it has a rotational symmetry of 360/n = 360/5 = 72 degrees. This means that the figure matches up with itself perfectly every 72 degrees of rotation. You are rotating this figure about the center." You have to try understanding rather than wait for someone to give you the answer, the square has an n of 4...\[\frac{ 360 }{ n } =? ~~~n = 4\]

OpenStudy (nightmarenight):

So it's 90? @iambatman

OpenStudy (nightmarenight):

Because 360 / 4 is 90.. @iambatman

OpenStudy (anonymous):

Yes